# Planning behavior in a recurrent neural network that plays Sokoban

### v1

@@ 2 Training the test subject
Guez et al. (2019) introduced the Deep Repeating ConvLSTM (DRC), whose core consists of $D$ convolutional LSTM layers with 32 channels and $3\times 3$ filters, each applied $N$ times per time step. Our $\text{DRC}(3,3)$ – or just DRC for brevity – has $1.29$M parameters. Before the LSTM core, two convolutional layers (without nonlinearity) encode the observation with $4\times 4$ filters.
The LSTM core uses $3\times 3$ convolutional filters, and a nonstandard tanh on the output gate (Jozefowicz et al., 2015). Unlike the original ConvLSTM (Shi et al., 2015), the input to each layer of a DRC consists of several concatenated components:
@@
  • To allow spatial information to travel fast in the ConvLSTM layers, we apply pool-and-inject by max- and mean-pooling the previous step’s hidden state. We linearly combine these values channel-wise before feeding them as an input to the next step.
@@
This is a convolutional residual neural network, also from Guez et al. (2019). It serves as a non-recurrent baseline that can only think during the forward pass (no ability to think for extra steps) but is nevertheless good at the game. The ResNet consists of 9 blocks, each with $4\times 4$ convolutional filters. The first two blocks have 32 channels, and the others have 64. Each block consists of a convolution, then two (relu, conv) sub-blocks, which each split off and are added back to the trunk. The ResNet has $3.07$M parameters.
@@
3 Quantitative behavior analysis
Figure 3: Proportion of medium-difficulty validation levels solved by DRC with $x$ time steps to think. The dotted line shows the ResNet’s solve proportion. The proportion of solved levels increases with thinking time, but only up to 6 steps.
3.1 Recurrent networks are better at Sokoban
After training both the recurrent $\text{DRC}(3,3)$ network and feed-forward ResNet for 2B environment transitions, the $\text{DRC}(3,3)$ network is able to solve $99.3\%$ of unfiltered test levels, while the ResNet solves only $97.9\%$. This gap is particularly notable given that the ResNet has more than double the parameters of the $\text{DRC}(3,3)$ network, so the ResNet has more model capacity and sequential thinking time. These results are consistent with Guez et al. (2019), see Appendix B.
The performance gap widens when considering just the hard levels, of which the $\text{DRC}(3,3)$ network solves $42.8\%$ and the ResNet $26.2\%$. Counterintuitively, we find that – controlling for performance on the unfiltered test set – recurrent networks are not comparatively better at harder levels than the ResNet. Figure 10 plots the performance of model checkpoints on the unfiltered test set against medium-difficulty and hard levels. We find that both model architectures follow the same curve: the recurrent $\text{DRC}(3,3)$ network is further to the right, but follows the same trend.
3.2 Thinking time improves performance
We find the performance of our trained $\text{DRC}(3,3)$ improves given additional thinking time, replicating Guez et al. (2019). We evaluate the DRC network on all medium-difficulty validation levels. To induce thinking time, we repeat the initial environment observation $n$ times while advancing the DRC hidden state, then let the DRC policy act normally. Figure 3 shows the DRC’s performance for $n\in\{0,1,2,4,6,8,10,12,16\}$ steps, alongside the ResNet baseline. The DRC’s success rate improves from $76.6\%$ to $81.3\%$ with more thinking steps, with peak performance at 6 steps. The effect ($4.7\%$ difference) is comparable to that observed by Guez et al. ($4.5\%$).
Figure 4: Average optimal solution length of levels grouped by the number of thinking steps at which the level is first solved. Each of the 50,000 medium-difficulty validation levels is assigned to a group. Optimal solution length is a proxy for difficulty, and we can see that levels which take longer to solve tend to be harder.
3.3 Planning emerges early in training
Figure 1 shows the performance of the DRC on the medium-difficulty validation levels by training step. We see the DRC rapidly acquires a strong planning effect by 70M environment steps. This effect strengthens on the hard levels as training progresses, but slowly weakens for medium-difficulty levels. This suggests that the training setup and loss function reinforce the planning effect for harder levels and disincentivize it for levels that are comparatively easier.
3.4 Do difficult levels benefit more from thinking?
The A* solver gives us two proxies for the difficulty of a level: the length of the optimal solution, and the number of nodes it had to expand to find a solution. Figure 4 shows a positive correlation between the A* optimal solution length ($y$-axis) and the number of DRC thinking steps $n$ needed to solve a level ($x$-axis). In other words, taking extra more thinking time helps solve harder levels.
On the other hand, Figure 11 (Appendix D) finds only a weakly positive correlation between the expanded nodes proxy and required thinking steps $n$, with results noisy for smaller $n$. We speculate that the ‘expanded nodes’ proxy metric is only weakly related to required $n$ due to the DRC having different heuristics than the one we use for A*, whereas the optimal solution length does not depend on heuristics.
Table 1: Levels at 6 thinking steps.
@@ 3.4 Do difficult levels benefit more from thinking?
3.5 Other consequences of DRC thinking time
While Figure 3 shows that the DRC can solve many more levels when given thinking steps at the beginning of an episode, in this section, we explore how the network’s behavior changes with thinking time, and if this has significant effects besides solving more levels.
Thinking makes the DRC “patient.”
One hypothesis is that, without the extra thinking steps, the DRC sometimes comes up with a greedy solution for some of the boxes and executes it right away. This puts some boxes into the wrong targets, and makes the puzzle unsolvable. Since the training signal penalizes every step it takes to solve the puzzle, it could be optimal for a bounded rational agent to avoid thinking and be “impatient” in this way: the agent may accumulate more return on average by being “impatient” and quickly solving most levels even if it sometimes errs by taking actions that make it impossible to solve other levels.
For $n$ thinking steps, Figure 5 plots the average number of steps (time-to-box) until each of 4 boxes gets on a target, in the order in which they do so. The results support the hypothesis: levels that are solved at 6 thinking steps but unsolved at 0 thinking steps show increased time-to-box for boxes 1-3 but decreased time for box 4 (Figure 5(b)). In contrast, averaged across all levels times go down for all boxes with increased thinking (Figure 5(a)).
Thinking often improves solutions for already-solved levels, but is sometimes harmful.
Table 1 shows that an additional 6 thinking steps solves 6.87% of previously unsolved levels – but at the cost of no longer being able to solve 2.23% of previously solved levels, for a total improved solution rate of 4.64%. In addition to improving the solution rate, additional thinking improves returns on average by shortening solution length in 18.98% of episodes, consistent with the decreased timesteps to place boxes shown in Figure 5(a). However, the effect is inconsistent: most episodes show no change in returns, with returns even decreasing in 5.26% of episodes. Overall, extra thinking steps are not always beneficial but, on balance, they let the DRC solve more levels or improve return on already solved levels.
Figure 5: Average time step to place each box $Bi$ on target for different thinking steps. (a) Averages across all levels where the box $Bi$ is placed on target by DRC with $n$ thinking steps. If box $Bi$ is not placed (on unsolved levels) we ignore the box, but we keep the other boxes in the level. (b) Averages for all levels solved by $6$ thinking steps but not solved by 00 thinking steps.
(a) file 0, level 18
[…]
(b) file 0, level 53
[…]
(c) file 0, level 153
[…]
Figure 6: Case studies of three medium-validation levels demonstrating different behaviors after 6 thinking steps. Colors are as in Figure 2. Boxes and targets are paired in upper- and lower-case letters respectively, such that the DRC’s best solution places boxes on targets in alphabetical order. Videos available at this https URL; see Section 4 for behavior descriptions. Levels solved faster obtain higher return as the agent incurs the per-step penalty fewer times.
4 Case studies: how does thinking change behavior?
To gain a deeper understanding of the DRC networks’ behavior, we examine an example from each of three categories from Table 1: a solved, previously unsolved level; a level on which the DRC improved its return with extra thinking (solved faster); and a level on which the return worsens (solved slower). We select an illustrative example of each category from the first ten examples by level index.
Thinking lets the DRC solve: Figure 6(a).
In the no-thinking condition, the DRC first pushes box $C$ one square to the right. It then goes back to push $A$ to $a$, but it is too late: it is now impossible to push box $B$ onto $b$. The network pushes $C$ onto $c$, then $D$ onto $d$ and then remains mostly still. In contrast, after thinking, the DRC pushes $A$ to $a$ first, which lets it solve the level.
Thinking speeds up solving: Figure 6(b).
In the no-thinking condition, the DRC spends many steps going back and forth before pushing any boxes. First it goes down to $y$, then up to $c$, then down onto $z$, back up to $y$ and to $z$ again. It then proceeds to solve the rest of the puzzle correctly: push box $A$ onto $a$, prepare box $B$ on $x$ and box $C$ where $B$ originally was, push in boxes $B,C$ and finally $D$. In the thinking condition, the DRC makes a beeline for $A$ and then plays the exact same solution.
Thinking slows down solving: Figure 6(c).
In the no-thinking condition, the DRC starts by pushing box $C$ into position $y$, then pushes boxes $A,B$. On the way back down, ther DRC pushes $C$ onto $c$ and finally $D$ onto $d$. In contrast, after thinking for a bit, the DRC goes the other way and starts by pushing $B$ onto $b$. The solution is the same after that: $A$, $C$, then $D$; but the NN has wasted time trekking back from $B$ to $A$.
4.1 Hypothesis: the DRC ‘paces’ to get thinking time
The second case study above found the DRC sometimes moves around in cycles without touching any box. One hypothesis for why this occurs is that the DRC is using the extra steps to come up with a plan to solve the episode. To test this, we can check whether cycles in the game state occur more frequently near the start of the episode, and whether deliberately giving the DRC thinking time makes it stop going in cycles.
Consistent with this hypothesis, Figure 7 (left) shows the majority of cycles start within the first 5 steps of the episode, and (right) forcing it to think for six steps makes about 75% of these cycles disappear.
The episode start is not the only time at which the DRC moves in cycles. Figure 9 shows that replacing an $n$-length cycle with $n$ thinking steps leads the NN to have the exact same behavior for at least 60% of levels, for at least 30 steps after the cycle. For context, the median solution length for train-unfiltered is exactly 30 steps. Additionally, in $82.39\%$ of cases, doing this prevents cycles from starting in the $n$ steps after the thinking-steps conclude (Appendix C).
Why does this behavior emerge during training? Thinking is useful for achieving higher return, so it should be reinforced. But it also has a cost, -0.1 per step, so it should be discouraged in easy levels that do not need the computation. We speculate that, as training advances and heuristics get tuned, the DRC needs to think in fewer levels, and it is better at knowing when it needs to pace. This would also explain why, after many training steps, the DRC network begins to benefit less from additional thinking steps at the start of the episode in Figure 1: the DRC finds levels easier and also knows to pace when thinking is needed.
Figure 7: Left: Histogram of cycle start times on the medium-difficulty validation levels. Right: Total number of cycles the agent takes in the first 5 steps across all episodes in medium-difficulty validation levels with $n$ initial thinking steps.
5 Related work
Interpretability of agents and world models.
Several works have attempted to find the mechanism by which a simple neural network does planning in mazes (Ivanitskiy et al., 2023; Mini et al., 2023), gridworlds (Bloom & Colognese, 2023), and graph search (Ivanitskiy et al., 2023). We believe the DRC we present is a clearer example of an agent than what these works focus on, and should be similarly possible to interpret.
Other works have found emergent world models in sequence-prediction (Li et al., 2023) and navigation (Wijmans et al., 2023) neural networks.
Goal misgeneralization and mesa-optimizers for alignment.
From the alignment perspective, AIs optimizing monomaniacally for a goal have long been a concern (e.g. Yudkowsky, 2006; Omohundro, 2008; see the preface of Russell, 2019). In a machine learning paradigm (Hubinger et al., 2019), the goal of the training system is not necessarily optimized; instead, the NN may optimize for a related or different goal (Di Langosco et al., 2022; Shah et al., 2022), or for no goal at all.
Chain-of-thought faithfulness.
Large language models (LLMs) use chain of thought, but are they faithful to it or do they think about their future actions in other ways (Lanham et al., 2023; Pfau et al., 2024)? One could hope that LLMs perform all long-term reasoning in plain English, allowing unintended human consequences to be easily monitored, as in Scheurer et al. (2023).
Reasoning neural network architectures.
Many papers try to enhance NN thinking by altering the training setup or architecture (Bansal et al., 2022; Graves, 2016; Chung et al., 2024).
Ethical treatment of AIs.
Do AIs deserve moral consideration? Schwitzgebel & Garza (2015) argue that very human-like AIs are conceivable and clearly deserve rights. Tomasik (2015) suggests that most AIs deserve at least a little consideration, like biological organisms of any species (Singer, 2004). But what does it mean to treat an AI ethically? Daswani & Leike (2015) argue that the way to measure pleasure and pain in a reinforcement learner is not by its absolute amount of return, but rather by the temporal difference (TD) error: the difference between its expectations and the actual return it obtained. If the internals of the NN have a potentially different objective (Hubinger et al., 2019; Di Langosco et al., 2022), then the TD error should come from a place other than the critic. This paper is an early step toward finding the learned-reward internal TD error, if it exists.
6 Conclusion
We have developed and open-sourced a model organism for understanding reasoning in neural networks. In addition, we contribute a detailed empirical analysis of this network’s behavior. We believe this work will prove useful to the interpretability, alignment, and ethical treatment of AIs.
First, we have trained a DRC recurrent network that plays Sokoban and benefits from additional thinking steps at test time, replicating Guez et al. (2019).
Second, we have shown that thinking for longer helps solve harder levels, and makes the DRC better at levels that require longer-sighted behavior. Without intervention, thinking sometimes takes the form of the DRC ‘pacing’ at the beginning or middle of the episode, in a way which can be substituted by repeating the same input. This suggests the network is deliberately using more computation by pacing.
Finally, we have shown that the training setup incentivizes the planning effect at the start. We find that planning is disincentivized later, but only for easier levels. We offer a hypothesis about why the training process disincentivizes the planning effect: the NN finds levels easier (needs less thinking), and also learns when to do the thinking it needs (via pacing).
7 Acknowledgements
We would like to thank ChengCheng Tan for editing this paper, Philip Quirke for project management, and the rest of the FAR team for general support. We also thank Maximilian Li, Alex F. Spies, and anonymous reviewers for improving the paper with their comments and edits. We are also very grateful for the many questions about their RL setup that Arthur Guez and Timothy Lillicrap answered, as well as inspiration from Stephen Chung and Scott Goodfriend.
Appendix A Training hyperparameters
All networks were trained with the same hyperparameters, which were tuned on a combination of the ResNet and the $\text{DRC}(3,3)$. These are almost exactly the same as Guez et al. (2019), allowing for taking the mean of the per-step loss instead of the sum.
Loss.
Gradient clipping and epsilon
The original IMPALA implementation, as well as Huang et al. (2023), sum the per-step losses. We instead average them for more predictability across batch sizes, so we had to scale down some parameters by a factor of $1/640$: Adam $\epsilon$, gradient norm for clipping, and L2 regularization).
Weight initialization.
We initialize the network with the Flax (Heek et al., 2023) default: normal weights truncated at 2 standard deviations and scaled to have standard deviation $\sqrt{1/\text{fan\_in}}$. Biases are initialized to 0. The forget gate of LSTMs has 1 added to it (Jozefowicz et al., 2015). We initialize the value and policy head weights with orthogonal vectors of norm 1. Surprisingly, this makes the variance of these unnormalized residual networks decently close to 1.
Adam optimizer.
L2 regularization.
In the training loss, we regularize the policy logits with L2 regularization with coefficient $1.5625\times 10^{-6}$. We regularize the actor and critic heads’ weights with L2 at coefficient $1.5625\times 10^{-8}$. We believe this has essentially no effect, but we left it in to more closely match Guez et al. (2019).
Software.
We base our IMPALA implementation on Cleanba (Huang et al., 2023). We implemented Sokoban in C++ using Envpool (Weng et al., 2022) for faster training, based on gym-sokoban (Schrader, 2018).
A.1 Number of training steps
[…]
Appendix B Learning curve comparison
It is difficult to fully replicate the results by Guez et al. (2019). Chung et al. (2024) propose an improved method for RL in planning-heavy domains. They employ the IMPALA $\text{DRC}(3,3)$ as a baseline and plot its performance in Chung et al. (2024, Figure 5). They plot two separate curves for $\text{DRC}(3,3)$: that from Guez et al. (2019), and a decent replicated baseline. The baseline is considerably slower to learn and peaks at lower performance.
We did not innovate in RL, so were able to spend more time on the replication. We compare our replication to Guez et al. (2019) in Figure 8, which shows that the learning curves for $\text{DRC}(3,3)$ and ResNet are compatible, but not the one for DRC(1,1). Our implementation also appears much less stable, with large error bars and large oscillations over time. We leave addressing that to future work.
The parameter counts (Table 2) are very different from what Guez et al. (2019) report. In private communication with the authors, we confirmed that our architecture has a comparable number of parameters, and some of the originally reported numbers are a typographical error.
Table 2: Parameter counts for each architecture.
Architecture Parameter count
$\text{DRC}(3,3)$ 1,285,125 (1.29M)
DRC(1,1) 987,525 (0.99M)
ResNet: 3,068,421 (3.07M)
Figure 8: Success rate for Test-unfiltered and Validation-medium levels vs. environment steps of training. Each architecture has 5 random seeds, the solid line is the pointwise median and the shaded area spans from the minimum to the maximum. The dotted lines are data for the performance of architectures extracted from the (Guez et al., 2019) PDF file. The values are slightly different from what Figures 1 and 3 report because they are calculated on a random sample of 1024 levels (24 levels are repeated for test-unfiltered).
Training Env Test Unfiltered Valid Medium
Steps ResNet DRC(3, 3) ResNet DRC(3, 3)
Success Return Success Return Success Return Success Return
100M 87.8 8.13 95.4 9.58 18.6 -6.59 47.9 -0.98
500M 93.1 9.24 97.9 10.21 39.7 -2.64 66.6 2.62
1B 95.4 9.75 99.2 10.47 50.0 -0.64 70.4 3.40
2B 97.9 10.29 99.3 10.52 59.4 1.16 76.6 4.52
Appendix C Additional cycles investigation
We run the DRC(3,3) on medium-validation levels and record where cycles in the game state happen. We prune out the redundant cycles that visit the same states as a given cycle. There are in total 13702 non-redundant cycles. For each cycle, if it is of length $N$, we replace it with $N$ steps of thinking, and measure what happens. In 86.74% of cases the DRC(3,3) does not immediately start going in cycles after the thinking time. That is, no cycle begins in the step after thinking time. In 82.39% of cases this persists for the next $N$ time steps: no cycles start in the $N$ steps after thinking.

### v2

@@ Appendix A Training the test subject
Guez et al. [2019] introduced the Deep Repeating ConvLSTM (DRC), whose core consists of $D$ convolutional LSTM layers with 32 channels and $3\times 3$ filters, each applied $N$ times per time step. Our $\text{DRC}(3,3)$ – or just DRC for brevity – has $1.29$M parameters. Before the LSTM core, two convolutional layers (without nonlinearity) encode the observation with $4\times 4$ filters.
The LSTM core uses $3\times 3$ convolutional filters, and a nonstandard tanh on the output gate [Jozefowicz et al., 2015]. Unlike the original ConvLSTM [Shi et al., 2015], the input to each layer of a DRC consists of several concatenated components:
@@
  • To allow spatial information to travel fast in the ConvLSTM layers, we apply pool-and-inject by max- and mean-pooling the previous step’s hidden state. We linearly combine these values channel-wise before feeding them as input to the next step.
@@
This is a convolutional residual neural network, also from Guez et al. [2019]. It serves as a non-recurrent baseline that can only think during the forward pass (no ability to think for extra steps) but is nevertheless good at the game. The ResNet consists of 9 blocks, each with $4\times 4$ convolutional filters. The first two blocks have 32 channels, and the others have 64. Each block consists of a convolution, followed by two (relu, conv) sub-blocks, each of which splits off and is added back to the trunk. The ResNet has $3.07$M parameters.
@@
A.1 Training hyperparameters
All networks were trained with the same hyperparameters, which were tuned on a combination of the ResNet and the $\text{DRC}(3,3)$. These are almost exactly the same as Guez et al. [2019], allowing for taking the mean of the per-step loss instead of the sum.
Time limits.
During training, we want to prevent strong time correlations between the returns, so the gradient steps are not correlated over time. For this reason, the time limit for each episode is uniformly random between 91 and 120 time steps.
Loss.
Gradient clipping and epsilon
The original IMPALA implementation, as well as Huang et al. [2023], sum the per-step losses. We instead average them for more predictability across batch sizes, so we had to scale down some parameters by a factor of $1/640$: Adam $\epsilon$, gradient norm for clipping, and L2 regularization).
Weight initialization.
We initialize the network with the Flax [Heek et al., 2023] default: normal weights truncated at 2 standard deviations and scaled to have standard deviation $\sqrt{1/\text{fan\_in}}$. Biases are initialized to 0. The forget gate of LSTMs has 1 added to it [Jozefowicz et al., 2015]. We initialize the value and policy head weights with orthogonal vectors of norm 1. Surprisingly, this makes the variance of these unnormalized residual networks decently close to 1.
Adam optimizer.
L2 regularization.
In the training loss, we regularize the policy logits with L2 regularization with coefficient $1.5625\times 10^{-6}$. We regularize the actor and critic heads’ weights with L2 at coefficient $1.5625\times 10^{-8}$. We believe this has essentially no effect, but we left it in to more closely match Guez et al. [2019].
Software.
We base our IMPALA implementation on Cleanba [Huang et al., 2023]. We implemented Sokoban in C++ using Envpool [Weng et al., 2022] for faster training, based on gym-sokoban [Schrader, 2018].
A.2 Number of training steps
[…]
A.3 Learning curve comparison
It is difficult to fully replicate the results by Guez et al. [2019]. Chung et al. [2024] propose an improved method for RL in planning-heavy domains. They employ the IMPALA $\text{DRC}(3,3)$ as a baseline and plot its performance in Chung et al. [2024, Figure 5]. They plot two separate curves for $\text{DRC}(3,3)$: that from Guez et al. [2019], and a decent replicated baseline. The baseline is considerably slower to learn and peaks at lower performance.
We did not innovate in RL, so were able to spend more time on the replication. We compare our replication to Guez et al. [2019] in Figure 7, which shows that the learning curves for $\text{DRC}(3,3)$ and ResNet are compatible, but not the one for DRC(1,1). Our implementation also appears much less stable, with large error bars and large oscillations over time. We leave addressing that to future work.
The parameter counts (Table 4) are very different from what Guez et al. [2019] report. In private communication with the authors, we confirmed that our architecture has a comparable number of parameters, and some of the originally reported numbers are a typographical error.
Table 4: Parameter counts for each architecture.
Architecture Parameter count
$\text{DRC}(3,3)$ 1,285,125 (1.29M)
$\text{DRC}(1,1)$ 987,525 (0.99M)
ResNet: 3,068,421 (3.07M)
Figure 7: Success rate for Test-unfiltered and Validation-medium levels vs. environment steps of training. Each architecture has 5 random seeds, the solid line is the pointwise median and the shaded area spans from the minimum to the maximum. The dotted lines are data for the performance of architectures extracted from the [Guez et al., 2019] PDF file. The values are slightly different from what Figures 5 and 3 report because they are calculated on a random sample of 1024 levels (24 levels are repeated for test-unfiltered, which only has 1000 levels).
Training Env Test Unfiltered Valid Medium
Steps ResNet DRC(3, 3) ResNet DRC(3, 3)
Success Return Success Return Success Return Success Return
100M 87.8 8.13 95.4 9.58 18.6 -6.59 47.9 -0.98
500M 93.1 9.24 97.9 10.21 39.7 -2.64 66.6 2.62
1B 95.4 9.75 99.2 10.47 50.0 -0.64 70.4 3.40
2B 97.9 10.29 99.3 10.52 59.4 1.16 76.6 4.52
Appendix B Generalizing the $\text{DRC}(3,3)$ to larger levels
We license the levels by Þorsteinsson [2009] as GPLv3, and make them available in the supplementary material. The performance of the $\text{DRC}(3,3)$ on each set of levels is in Table 6. The sets with higher scores are also those which humans find easier: for example, “Dimitri & Yorick” was made for children by Jacques Duthen and consists entirely of relatively small levels (largest is $12\times 10$) with at most 5 boxes. The sets where the $\text{DRC}(3,3)$ solves nothing are also difficult for the authors of this paper.
We tried giving the $\text{DRC}(3,3)$ from 2 to 128 extra thinking steps for these levels, incrementing in powers of two. For most sets, we find some benefit to 2-4 thinking steps, but no more. The sole exception is XSokoban, which contains one level (Figure 2, right) which requires 128 steps of thinking to solve.
We encourage the reader to go to the website by Þorsteinsson [2009] and try solving the levels.
(a) Microban, level 105
(b) Microban, level 144
(c) Mas Sasquatch, level 15
Figure 8: Thinking steps are not always useful: For level (a), the network succeeds with 0 to 16 thinking steps, but fails with 32 or more thinking steps. Levels (b) and (c) are too difficult for the $\text{DRC}(3,3)$, and it always fails on them independent of thinking steps, though it comes up with decent partial solutions. All levels from Þorsteinsson [2009].
(a) Sokoban Jr. 2, level 17
(b) Microban, level 145
(c) XSokoban, level 30
Figure 9: Some levels require thinking steps for success: For level (a), the network always succeeds. For level (b), it succeeds at 32 or more thinking steps. For level (c), it needs 128 thinking steps to succeed. All levels from Þorsteinsson [2009].
Table 6: Performance of the $\text{DRC}(3,3)$ on each set of levels by [Þorsteinsson, 2009]. The “Max solved” columns represent the proportion of levels solved at the number of steps in the “at steps” column, which is the highest solved proportion for each number of thinking steps tried.
All levels Levels larger than $10\times 10$
Level collection # solved Max slv. max at # solved Max slv. max at
Dimitri & Yorick 61 83.6% 88.5% 16 0
Sokoban Jr. 1 60 80.0% 81.7% 2 19 73.7% 73.7% 0
Howard’s 3rd set 40 60.0% 62.5% 4 1 0.0% 0.0% 0
Simple sokoban 61 54.1% 60.7% 16 51 47.1% 54.9% 16
Sokoban Jr. 2 54 48.1% 50.0% 2 40 45.0% 47.5% 2
Microban 155 23.9% 24.5% 4 17 5.9% 11.8% 2
Deluxe 55 21.8% 23.6% 2 1 0.0% 0.0% 0
Sokogen 990602 78 20.5% 23.1% 8 0
Yoshio Automatic 52 13.5% 17.3% 2 0
Sasquatch III 16 6.2% 6.2% 0 8 0.0% 0.0% 0
Howard’s 1st set 100 5.0% 6.0% 2 54 0.0% 0.0% 0
Howard’s 2nd set 40 5.0% 10.0% 4 22 0.0% 0.0% 0
Microcosmos 40 5.0% 7.5% 4 0
Still more levels 35 2.9% 2.9% 0 34 2.9% 2.9% 0
Sasquatch IV 36 2.8% 2.8% 0 20 0.0% 0.0% 0
Xsokoban 40 2.5% 5.0% 128 39 2.6% 5.1% 128
Sasquatch 49 2.0% 4.1% 4 39 0.0% 2.6% 4
David Holland 1 10 0.0% 0.0% 0 5 0.0% 0.0% 0
David Holland 2 10 0.0% 0.0% 0 9 0.0% 0.0% 0
Howard’s 4th set 32 0.0% 0.0% 0 30 0.0% 0.0% 0
Mas Sasquatch 50 0.0% 0.0% 0 43 0.0% 0.0% 0
Nabokosmos 40 0.0% 0.0% 0 0
Sokoban 50 0.0% 0.0% 0 48 0.0% 0.0% 0
Appendix C Case Studies
(a) file 0, level 18
[…]
(b) file 0, level 53
[…]
(c) file 0, level 153
[…]
Figure 10: Case studies of three medium-validation levels demonstrating different behaviors after 6 thinking steps. Colors are as in Figure 12. Boxes and targets are paired in upper- and lower-case letters respectively, such that the DRC’s best solution places boxes on targets in alphabetical order. Videos available at this https URL. Levels solved faster obtain higher return as the agent incurs the per-step penalty fewer times. Note that the letters used are for reference and not a property of Sokoban.
Case Study: Thinking makes some levels solvable Figure 10(a).
For example, thinking lets the DRC solve Figure 10(a). In the no-thinking condition, the DRC first pushes box $C$ one square to the right. It then goes back to push $A$ to $a$, but it has now become impossible to push box $B$ onto $b$. In contrast, after thinking, the DRC pushes $A$ to $a$ first, which lets it solve the level.
Case Study: Thinking speeds up solving Figure 10(b).
Figure 10(b) illustrates a scenario where the agent happens. In the no-thinking condition, the DRC takes many steps back and forth before pushing any boxes. First it goes down to $y$, up to $c$, then down onto $z$, back up to $y$ and to $z$ again. It then proceeds to solve the rest of the puzzle: push box $A$ onto $a$, prepare box $B$ on $x$ and box $C$ where $B$ originally was, push in boxes $B,C$ and finally $D$. In the thinking condition, the DRC makes a beeline for $A$ and then plays the same solution.
Case Study: Thinking slows downFigure 10(c).
Figure 10(c) illustrates a scenario where thinking time results in a slower solution. In the no thinking condition, the DRC starts by pushing box C into position y, then pushes boxes A, then B into place. On the way back down, the DRC pushes C onto c and finally D onto d. In contrast, in the thinking condition, the DRC goes the other way and starts by pushing B onto b. The subsequent solution (A, C, then D) is the same, but the DRC has wasted time trekking back from B to A.
Appendix D Probe Training
Probe architecture
We train multiple different linear probes on the hidden states $h$ and cell states $c$ activations of the network. The probes are trained either on the states of the individual layers or by concatenating the states of all the three layers. We collect the activations of the model by letting it play through the hard levels. The training set for the probes are constructed by taking a random sample of cached levels and including all the timesteps in the level excluding the first five timesteps. We exclude these intial steps as we believe they can be noisy as the network comes up with the stategy to solve the levels.
The probes are trained with logistic regression with L1 decay using the Scikit-Learn library. We do a grid-search on the learning rate and the L1 weight decay to select the probe that has the highest F1-score on a validation set.
For multi-class probe targets, each potential output class $l$ is treated like a different data point. That is: for each label $l$, we consider a prediction positive if the highest probe logit is for label $l$, and negative otherwise. We consider a data point positive iff the true label is label $l$. Then, we compute the confusion matrix and the F1 score from the $n\cdot l$ data points.
Appendix E Details about Sparse Autoencoders (SAEs)
Figure 11: Visualization of some interpretable features from the SAE of last layer. These features also appear monosemantically in the channels. The precision, recall, and F1 score for the features are reported in Table 10.
In order to search for more monosemantic and interpretable features in the network, we train sparse-autoencoders (SAEs) [Huben et al., 2023, Bricken et al., 2023] on the individual squares in the $h$ hidden state of the network consisting of 32 neurons. Thus, we get a $10\times 10$ visualization for each SAE feature as shown in Figure 11. We use the top-k activation function [Gao et al., 2024, Tamkin et al., 2024] for the SAE as it is currently the state-of-the-art method to train SAEs and is easier to work with as it directly sets the $L_{0}$ norm of the SAE activations. The hyperparameter search space for training the SAE is provided in Table 7. We train separate SAEs for each layer with the specified hyperparameters and pick the one that achieves greater than $90\%$ explained variance while having interpretable features assessed through manual visual inspection. We release these trained SAEs and probes in the same huggingface repo as our trained DRC networks.  [^8] https://huggingface.co/AlignmentResearch/learned-planner
We provide examples of some of the interpretable features from an SAE trained on the last layer with $k=8$ in Table 8 with corresponding visualization in Figure 11. For the “Target”, “Unsolved”, and “Solved” concepts (Table 10), we observed cases where the SAE feature is offset from the ground truth by 1 square in either horizontal or vertical directions. The level setup includes a permanent one-square outer-edge wall, so this offset never results in an out-of-bounds issue. We evaluated these potential “Offset” variants for the Target, Unsolved, and Solved concepts.
All of the features in the SAE that we find to be interpretable are already embedded in individual channels. These channels are either as monosemantic as the SAE features, or in some cases more monosemantic as measured by the F1 score. Table 3 reports the precision, recall, and F1 scores for the action features as evaluated through channels, SAE neurons, and linear probe trained against the ground truth action predictions. Table 10 reports the same scores for additional interpretable features. The SAE action features have lower F1 scores than channels by a margin of $5.9\%$ on average. On the other hand, the linear probes trained across all channels of the hidden state have similar F1 scores as the specified channels indicating that the channels are monosemantic and cannot be improved upon by combining various channels using a linear probe.
Table 7: Hyperparameter search space for training SAE
Hyperparameter Search Space
$k$$\{4,8,12,16\}$
learning rate $\{1e-5,5e-5,1e-4,5e-4,1e-3\}$
expansion factor $\{16,32,64\}$
Table 8: SAE Feature Concepts
Concept Description
Target The 4 target squares (static)
Unsolved Targets and boxes that aren’t solved
Solved Solved target squares with a box on them
Agent Up The agent will move Up next step
Agent Down The agent will move Down next step
Agent Left The agent will move Left next step
Agent Right The agent will move Right next step
Table 9: Breakdown of levels by category at 6 thinking steps.
@@ Appendix E Details about Sparse Autoencoders (SAEs)
Table 10: Scores for SAE and Channel features
Concept Offset $(dy,dx)$ Channel SAE Feature
Number Prec Rec F1 Number Prec Rec F1
Target (1, 0) L3C17 97.8 97.7 97.8 L3F278 97.8 98.1 98.0
Unsolved targets and boxes (0, 0) L3C7 94.9 90.8 92.8 L3F212 95.3 86.6 90.7
Solved targets (0, 0) -L3C7 91.6 94.6 93.0 L3F179 91.7 91.5 91.6
Appendix F Additional quantitative behavior figures and tables
Figure 13: Number of thinking steps required to solve the level vs. number of nodes A* needs to expand to solve it. The trend is increasing at the end but very unclear, indicating different heuristics used by the DRC and A*.